. If nominal GDP in a particular year is $770 billion while real GDP is $700 billion, we know that since the base year
Answer #1

GDP Deflator = (Nominal GDP / Real GDP) x 100 = (770 / 700) x 100 = 110

Since inflation rate is measured by % Change in GDP Deflator,

Inflation rate = (110/100) - 1 = 1.1 - 1 = 0.1 = 10%

Since the base year, Inflation rate increased by 10%.